# Categorical Flow Matching on Statistical Manifolds

### v1

@@ B.2 Bits-Per-Character (BPC) Calculation
We have demonstrated that the NLL can reach negative infinity. In contrast, in our definition of BPC with the cross-entropy loss $-\log_{2}p(\delta|\mu)$, it is always non-negative and consistent with previous autoregressive language models. However, as our flow-based model is not autoregressive, we instead follow [50] to calculate the pseudo-NLL/BPC for comparison. The pseudo-log-likelihood of a $D$-dimensional data is defined as
$\frac{1}{D}\sum_{d=1}^{D}\log p(x^{d}|x^{-d})$ (42)
where $x^{-d}=\{x^{k},1\leq k\leq D,k\neq d\}$ represents the data without the specific dimension index $d$. BPC can be similarly defined as $-\frac{1}{d}\sum_{d=1}^{D}\log_{2}p(x^{d}|x^{-d})$. In our evaluation with Text8, we always pick the middle token with a window of 128 known tokens before and 127 known tokens after. The model tries to recover the missing token given the conditional context. As our model only accepts one timestep for all data dimensions, instead of fixing the context, we still uniformly randomly sample all the initial noises from the simplex but substitute the corresponding vector fields with the ground truth ones to make sure the context tokens converge to the ground truth one-hot distributions.
@@ C.1 Model Parameterization
The projection guarantees that the final prediction lies on the corresponding tangent space, and the data points will stay on the manifold during Euler sampling. We will always assume the projection is performed in the following context of using $v_{t}$. The training stage of SFM is described in Alg.2. The time complexity of our SFM framework is dominated by the underlying vector field predictor with little overhead. Each model for binarized MNIST and promoter design was trained on a single NVIDIA A100 GPU for 6-10 hours. Each model for Text8 was trained on 8 NVIDIA A100 GPUs for about 7 days. We will further describe the encoders for each generation task in the following subsections.
@@
6:     Apply the diffeomorphism in Eq.5 to obtain $x_{0}=\pi(\mu_{0}),x_{1}=\pi(\mu_{1})$.
@@ C.2 Model Sampling
2:Apply the diffeomorphism in Eq.5 to obtain $x_{0}=\pi(\mu_{0})$.
@@
C.3 Swill Roll
The Swill roll on the 2-simplex is generated by normalizing the span of the 2-dimensional Swill roll to $[0+\varepsilon,1-\varepsilon]$ where $\varepsilon$ is a small margin to make sure no point lies on the boundary. The third dimensional is automatically obtained as $\mu_{3}=1-\mu_{1}-\mu_{2}$. We fixed a random seed and generated 1000 samples for a full batch training. The vector field predictor is based on simple multi-layer perceptions (MLPs) and can be described as
@@ C.4 Binarized MNIST
For the quantitative evaluation of the generated samples, we calculated the FID score for 1000 generated samples for each model. The statistics for the binarized MNIST dataset were calculated on the whole training set via the pre-trained InceptionV3 [57] model. The NLLs reported in the main text were calculated on a fixed subset with 1000 samples of the test set using the Dopri5 ODE solver.
@@ C.5 Text8
We followed previous work [23, 6] to use a fixed split of 90M/5M/5M with a fixed sequence length of 256. We used a 12-layer transformer-based model modified from [47] with an additional additive time embedding. The resultant model has 85.1M trainable parameters. Noticeably, our model size is smaller than [23] which used a full 24-layer model, but is similar to [37]. The models were trained for 1M iterations with a batch size of 512 (approximately 23 epochs), an initial learning rate of $10^{-4}$, and an exponential moving average (EMA) decay rate of 0.9999. We further used 1000 iterations as the linear warmup of the learning rate and used a decay rate of 0.8 with a patience of 3 and a validation interval of 2000 training iterations. The model snapshot with the lowest validation loss was saved for evaluation. BPC and NLL were calculated on the first 10k characters of the test set with 256 Euler steps, and generated texts were obtained using either 256 Euler steps or the Dopri5 ODE solver (see Appendix D.2).
@@ Appendix D Additional Result
In this section, we provide additional results of ablation studies as well as generated samples for each task.
D.1 Ablation Study
We provide ablation studies of the effect of different sampling methods, different sampling steps for the Euler methods, and different $t_{\text{max}}\approx 1$ for NLL calculation. The results are provided in Tab.5. The effect of the variational timestep $t_{\text{max}}$ matched the results from [7], which stated that a closer timestep to the target data would decrease the NLL. The reasons we choose $t=0.995$ instead of higher values are: 1) we noticed numerical stability issues at a timestep very close to the target, 2) we want a fair comparison with [7] which used an equivalent timestep of 0.9975, and 3) as long as we use a same timestep, the results are comparable within. We also found that NLL increased with the number of Euler sampling steps. We hypothesize that it is due to the large divergence change at $t\approx 1$ and naive Euler steps tend to overestimate its contribution. Nonetheless, the Euler method with more than 300 steps gave a similar NLL as the ODE solver. We further provide additional results of the FID scores on different sample methods in Tab.5 in which we used $t=0.995$ and 300 Euler steps. The final performance was quite similar between these two sampling methods.
@@ D.1 Ablation Study
Table 5: FID for different sampling methods with $t_{\text{max}}=0.995$. For the Euler method, we used 300 steps.
@@
ODE SFM w/ OT -1.687 $\pm$ 0.020
SFM w/o OT -1.631 $\pm$ 0.027
LinearFM 0.622 $\pm$ 0.022
Euler SFM w/ OT -1.689 $\pm$ 0.031
SFM w/o OT -1.653 $\pm$ 0.028
LinearFM 0.499 $\pm$ 0.022
D.2 Generated Samples
@@
The ground truth images and generated samples of the binarized MNIST dataset using various models and different sampling settings are demonstrated in Fig.4. The lower row demonstrates generated samples with different numbers of Euler steps. Generated texts on the Text8 dataset using different models and sampling methods are demonstrated below.
SFM w/ OT, ODE
a_gull_a_clock_as_sing_cowbreaks_diseases_in_forms_the_eilewes_of_clocks_contained_four_two_scarters_to_be_cursed_norse_the_hygg
y_alnanking_wade_of_joul_veine_in_one_ovias_sn_gl_cz_asi_evic_hal_art_diseashe_the_final_cangent_oedesis_spurl_the_last_one_surf
_cut_much_time_tow_through_the_wine_the_merlury_and_sore_characterography_is_lumaniua_nuzgen_by_khira_gormmaster_johk_herself_ha
s_nuquen_heads_some_conguees_following_row_the_living_core_coy_or_lanon_for_the_processions_and_workspandiios_for_three_e_conque
_was_lirelying_the_practice_of_finally_assista_he_was_one_of_the_superl_writings_though_using_the_qurdish_sieno_ardio_woo_daagul
itze_joy_to_pann_him_arrived_in_the_muslim_s_liorum_adam_and_were_ardius_who_was_to_have_not_thing_an_author_and_the_descent_hot
SFM w/o OT, ODE
running_of_experienting_and_preserved_in_to_them_no_longion_to_skin_this_is_not_serving_the_s_ready_pvorq_t_in_pchwadenaiosence_
bastiforces_in_wii_equal_partful_coward_dringetatt_is_often_devotes_to_devine_castles_faitst_gat_i_m_chilishka_fringelfat_to_the
s_help_to_lead_is_rich_to_his_court_the_acts_of_a_brateful_sharp_act_and_were_neploed_in_guilt_money_is_lost_recents_was_depende
nt_on_sweptumppichard_depeazed_n_using_ly_to_ruinto_war_the_kinbalt_rativation_that_will_snow_beploed_many_scritics_were_not_oft
_god_eight_one_doe_eight_german_embinor_eranger_and_the_fondam_s_andersen_diemstrop_filof_eight_one_silor_base_neget_gunshting_o
ther_pethors_in_support_the_texas_series_of_bright_pubtle_snap_czech_dogagova_s_thiru_buabing_aviator_space_k_ine_i_line_lines_a
LinearFM, ODE
_nine_three_zero_directions_requintly_the_post_wio_pranixenel_is_if_josalsted_but_shark_for_the_acchust_zoo_in_featural_acgule_o
f_micvilland_fame_there_two_nixed_by_for_the_aideadeqs_never_re_aine_is_mao_bun_jacubs_miles_for_ceatn_that_lost_not_roduitalist
form_lateuf_gh_movieg_thingporeefictine_slill_actral_the_dicav_siguraded_or_the_dorlas_oodi_ledic_missimines_a_consumts_tlaw_ref
arndy_by_hross_the_pous_by_darsick_murejfant_folloals_alenic_pnsemide_it_the_gremustate_fictional_and_tricity_to_to_the_obina_te
n_coan_in_treaty_of_areas_two_info_z_and_a_co_putu_for_the_rumo_stillia_defensive_productia__dit_one_zero_six_two_t_pe_itay_cdon
_ca_seven_as_used_as_mook_rejeired_space_foodrestions_cr_see_that_they_weure_some_pute_a_in_allime_amtc_s_oft_call_authought_pen
SFM w/ OT, Euler 256
by_keith_pang_ways_go_english_mouyr_hoch_in_one_nine_five_zero_th_ow_people_hugh_era_pizware_constitution_of_its_one_nine_five_o
ne_in_this_midway_co_english_anarchist_formed_clitue_these_america_produced_a_view_of_the_two_baggoots_and_america_mf_ayain_spen
icies_the_government_it_shot_to_replace_a_work_chagping_is_r_given_only_to_puddue_caparty_adminisment_in_two_zero_zero_three_au_
feez_giorrella_staffie_stewars_alfree_a_bum_often_boalhieved_by_significan_performances_of_axotas_a_in_the_world_wars_a_limit_fo
s_parented_by_two_ecll_which_repromite_s_tvhq_which_have_been_card_they_are_their_protrant_scorring_us_two_zero_million_raiters_
attempt_to_swing_a_pent_chatizen_of_the_tane_the_west_resattled_this_partement_are_anting_pasent_with_conse_x_dot_n_les_carling_
SFM w/o OT, Euler 256
art_of_the_dad_celsu_later_history_evea_s_d_gault_nentoci_lan_than_jans_grais_in_tood_s_daily_hobred_panes_than_was_recorded_in_
the_plot_of_dwaneke_s_elee_fate_carter_and_fauda_raylnell_mate_of_eveluaea_it_was_magnitheis_bereal_yearso_s_description_of_his_
_line_is_barer_edement_breeleys_lords_one_nine_seven_seven_gesadius_one_nine_eight_four_at_mike_cornelly_fonohre_lindon_pantheri
ngs_one_nine_eight_seven_one_line_the_cedenasipcs_the_ground_of_inci_are_gaurie_joining_the_story_and_tried_with_caway_inci_are_
f_strines_bose_such_as_throoptero_watermade_scuumenis_roche_and_streubs_the_unrbridted_apollophagia_i_two_five_zero_and_attends_
to_one_of_the_samkath_uses_maintaining_the_mosqui_tetonmeters_have_also_published_negau_in_place_ho_coubt_that_the_phoroachorido
LinearFM, Euler 256
h_jo_leek_the_form_buu_and_a_art_in_the_buel_he_bbu_bluiated_bpc_relations_of_debated_issued_that_age_denintenede_to_advers_adge
n_wmather_alconsensifia_cabip_lod_pyer_vubibax_which_s_vesskn_comesignearclij_ly_was_flee_down_becauid_the_pawer_of_cantix_are_s
tormut_of_duelibp_cas_unquests_pet_trokety_cal_totat_taiutext_two_logy_their_cylpheighfoding_in_par_has_sor_amale_optets_of_spyw
ork_for_sleampter_re_davi_cadiual_age_gestricational_gaie_the_cases_autharity_ouly_pop_radmitius__mamfare_the_diatuary_s_warking
w_cerip_mike_the_devisoll_rite_stringri_cometdral_king_hort_eitry_vix_orth_other_centogyika_venin_weads_nine_stan_shiup_extm_ziw
_xtermal_yy_tocharsical_lyvi_in_one_eighteight_peri_e_entoguian_drag_wham_paris_sin_res_nrigidal_ort_centric_gas_live_perby_the_

### v2

@@ B.2 Bits-Per-Character (BPC) Calculation
We have demonstrated that the NLL can reach negative infinity. In contrast, the common definition of BPC with the cross-entropy loss $-\log_{2}p(\delta|\mu)$ is always non-negative and consistent with previous autoregressive language models [61]. However, as our flow-based model is not autoregressive, we instead follow the variational formulation in [51] to calculate the alternative NLL and BPC for comparison. The ELBO in [51] can be calculated as
[…]
where $\mu_{t}=\exp_{\mu_{0}}(t\log_{\mu_{0}}\mu_{1}),\mu_{0}\sim p_{0}(\mu)$ denotes the interpolation along the geodesic and $\hat{\mu}_{1}(\mu_{t})$ denotes the learned model’s prediction for $t=1$ given the current noise data $\mu_{t}$. This can be efficiently done by one-step prediction with the learned vector field $v$ as
$\hat{\mu}_{1}(\mu_{t})=\exp_{\mu_{t}}\left((1-t)v_{t}(\mu_{t})\right).$ (44)
The standard Euclidean inner product is used in Eq.(43) so it can be understood as a weighted cross-entropy loss. Note that the integral in Eq.(43) has a singularity at $t=1$, making it numerically unstable to estimate using ODE solvers. Instead, we follow [51] to apply the change of variable $s=-\log(1-t)$ to reformulate the ELBO as
[…]
where $t=1-e^{-s}$. Note that for large $s$, the corresponding $t$ is very close to 1, so the integrand is very close to zero. Indeed, $s=10$ corresponds to $1-t<5\times 10^{-5}$, so we simply set the upper limit to 10 and used the Dopri5 solver to numerically estimate the ELBO. BPC can be then calculated as $\mathcal{L}_{\text{ELBO}}/\log 2$. This formulation also shares a similar form as the other ELBOs derived in [6, 12], and is thus comparable to most existing models.
Eq.(43) was directly optimized in Markov chain-based methods like D3PM [6], MultiFlow [12], and SEDD [38], and also autoregressive language models. In contrast, our SFM (and LinearFM) follows the continuous normalizing flow setup in [37], which enables the exact likelihood calculation instead of ELBOs. Therefore, trained on the alternative objective of minimizing the Riemannian vector field norm in Eq.(8), SFM did not directly try to minimize such an ELBO. Despite such an unfavorable evaluation compared to other baselines, SFM was still able to achieve the second-best BPC, as demonstrated in Tab.3.
B.3 GPT-J-6B Likelihood
GPT-J-6B [66] is a transformer-based large language model with 6B trainable parameters. We follow the pipeline in [12] to first tokenize the generated text using the provided tokenizer with GPT-J-6B. The GPT-J-6B NLL is then calculated based on the predicted logits on the token level using the pre-trained model as:
$\mathcal{L}_{\text{GPT}}=-\frac{1}{K}\sum_{k=1}^{K}\log p\left(w_{k}|w_{1:k-1}\right)$ (46)
where $w$ are tokens and $K$ is the number of tokens. Similarly, the entropy is calculated on the empirical distribution of the tokens based on a large number of generated texts. We noted that, as GPT-J-6B was not pre-trained on Text8, its NLL does not necessarily reflect the true data distribution in Text8, as we will demonstrate more concretely in the evaluation plot in Appendix D.1 and generated samples in Appendix D.3.
@@ C.1 Model Parameterization
The projection guarantees that the final prediction lies on the corresponding tangent space, and the data points will stay on the manifold during Euler sampling. We will always assume the projection is performed in the following context of using $v_{t}$. The training stage of SFM is described in Alg.2. The time complexity of our SFM framework is dominated by the underlying vector field predictor with little overhead. Each model for binarized MNIST and promoter design was trained on a single 80GB NVIDIA A100 GPU for 6-10 hours. Each model for Text8 was trained on four 80GB NVIDIA A100 GPUs for about 7 days. We will further describe the encoders for each generation task in the following subsections.
@@
6:     Apply the diffeomorphism in Eq.(5) to obtain $x_{0}=\pi(\mu_{0}),x_{1}=\pi(\mu_{1})$.
@@ C.2 Model Sampling
2:Apply the diffeomorphism in Eq.(5) to obtain $x_{0}=\pi(\mu_{0})$.
@@
C.3 Swiss Roll
The Swiss roll on the 2-simplex is generated by normalizing the span of the 2-dimensional Swiss roll to $[0+\varepsilon,1-\varepsilon]$ where $\varepsilon$ is a small margin to make sure no point lies on the boundary. The third dimensional is automatically obtained as $\mu_{3}=1-\mu_{1}-\mu_{2}$. We fixed a random seed and generated 1000 samples for a full batch training. The vector field predictor is based on simple multi-layer perceptions (MLPs) and can be described as
@@ C.4 Binarized MNIST
For the quantitative evaluation of the generated samples, we calculated the FID score for 1000 generated samples for each model. The statistics for the binarized MNIST dataset were calculated on the whole training set via the pre-trained InceptionV3 [60] model. The NLLs reported in the main text were calculated on a fixed subset with 1000 samples of the test set using the Dopri5 ODE solver. See Appendix D.3 for generated images.
@@ C.5 Text8
We followed previous work [23, 6] to use a fixed split of 90M/5M/5M with a fixed sequence length of 256. We used a 12-layer diffusion transformer (DiT) [48] based predictor, similar to the one used in [38]. The resultant model has 92.4M trainable parameters. Noticeably, our model size is smaller than [23] which used a full 24-layer model, but is similar to [38]. The models were trained for a total number of 3M iterations with a batch size of 512 per GPU (approximately 16 epochs), an initial learning rate of $10^{-4}$, and an exponential moving average (EMA) decay rate of 0.9999. We further used 1000 iterations as the linear warmup of the learning rate and used a decay rate of 0.8 with a patience of 3 and a validation interval of 2000 training iterations. The model snapshot with the lowest validation loss was saved for evaluation. BPCs were calculated on the first 4k sequences of length 256 of the test set (about 1M characters) with the Dopri5 solver as described in Appendix B.2, and generated texts were also obtained using the Dopri5 ODE solver. GPT-J-6B NLL and the generation token entropy were calculated based on 4k generations of length 256 (about 1M characters). See Appendix D.3 for generated texts.
@@ Appendix D Additional Result
In this section, we provide additional results of the evaluation on Text8, additional ablation studies, and generated samples for each task.
D.1 Additional Result on Text8
Figure 4: GPT-J-6B NLL versus sample entropy. For MultiFlow, D3PM, and autoregressive, the curve represents different logit temperatures from 0.5 to 1. Baseline data are from [12].
We further follow [12] (MultiFlow) to calculate the NLL using GPT-J-6B [66], a pre-trained autoregressive large language model. As such an NLL can be easily fooled by repeating high-frequency words, we also follow [12] to report additional token entropy, for which a closer entropy to data is preferred. The GPT-J-6B NLL and the token entropy are plotted in Fig.4, with the raw baseline data and the permission to reproduce provided by the authors of [12]. An additional random baseline that generates each character independently based on the letter frequency was provided. The ground truth data NLL and entropy are also included as the dotted horizontal and vertical lines. They demonstrate the best results any model can possibly achieve. Our SFM tended to better capture the diversity of the text corpus with the best entropy. Generated texts are provided in Appendix D.3.
As GPT-J-6B was not trained on Text8, we noted a caveat that its NLL may not necessarily reflect the Text8 distribution fitness. From the MultiFlow results, such an NLL can be made artificially low by duplicating high-frequency words, e.g., repeated numbers with little semantic meaning as in the low-temperature MultiFlow generations (see Appendix D.3 for examples). We also noted that such an NLL can be easily fooled by randomly generated strings based on letter frequency. Additionally, low-temperature MultiFlow variants achieved lower NLLs than the ground truth data, making this metric less credible. We noted the huge impact of temperature on MultiFlow as it generated more repeated numbers with lower temperatures. In contrast, SFM achieved perceptually similar or even better results with more diversity.
D.2 Ablation Study
We provide ablation studies of the effect of different sampling methods, different sampling steps for the Euler methods, and different $t_{\text{max}}\approx 1$ for NLL calculation. The results are provided in Tab.5. The effect of the variational timestep $t_{\text{max}}$ matched the results from [7], which stated that a closer timestep to the target data would decrease the NLL. The reasons we choose $t=0.995$ instead of higher values are: 1) we noticed numerical stability issues at a timestep very close to the target, 2) we want a fair comparison with [7] which used an equivalent timestep of 0.9975, and 3) as long as we use a same timestep, the results are comparable within. We also found that NLL increased with the number of Euler sampling steps. We hypothesize that it is due to the large divergence change at $t\approx 1$ and naive Euler steps tend to overestimate its contribution. Nonetheless, the Euler method with more than 300 steps gave a similar NLL as the ODE solver. We further provide additional results of the NLLs and FID scores on different sample methods in Tab.5 in which we used $t=0.995$ and 300 Euler steps. The final performance was quite similar between these two sampling methods.
@@ D.2 Ablation Study
Euler 300 0.9995 -2.545 $\pm$ 0.029
0.999 -2.442 $\pm$ 0.018
0.995 -1.689 $\pm$ 0.031
0.99 -1.019 $\pm$ 0.032
100 0.995 -1.871 $\pm$ 0.020
500 -1.677 $\pm$ 0.022
1000 -1.647 $\pm$ 0.024
ODE - 0.995 -1.687 $\pm$ 0.020
Table 5: NLL and FID for different sampling methods with $t_{\text{max}}=0.995$. For the Euler method, we used 300 Euler steps.
@@
ODE SFM w/ OT -1.687 $\pm$ 0.020 4.62
SFM w/o OT -1.631 $\pm$ 0.027 5.15
LinearFM 0.622 $\pm$ 0.022 5.91
Euler SFM w/ OT -1.689 $\pm$ 0.031 5.00
SFM w/o OT -1.653 $\pm$ 0.028 4.86
LinearFM 0.499 $\pm$ 0.022 6.47
D.3 Generated Samples
@@
The ground truth images and generated samples of the binarized MNIST dataset using various models and different sampling settings are demonstrated in Fig.5. The lower row demonstrates generated samples with different numbers of Euler steps.
For the Text8 dataset, generated texts, NLL, entropy, and per sample NLL evaluated by GPT-J-6B using different models and sampling methods are demonstrated in Tab.6, with one sample with an NLL chosen around the average, one above the average, and one below the average for each model to provide a more comprehensive demonstration of the generation. The MultiFlow with different sampling temperatures used the model checkpoint in [12]. As SFM directly samples categorical probabilities instead of making multiple discrete jumps with logits, it is incompatible with temperature-tuning.
Table 6: Generated samples and GPT-J-6B NLLs. For each model, the average NLL and entropy calculated on 4k generations are also provided. The NLL marked for each generation is the sample NLL, as evaluated by GPT-J-6B.
SFM w/ OT, ODE, NLL: 6.762, Entropy: 7.340
zero_zero_zero_more_as_well_as_the_needed_of_all_of_it_church_the_country_s_higner_upcoming_bank_the_country_comment_on_quebec_e
dits_includes_the_account_of_diego_hyle_ciaspare_coes_tain_three_zero_seven_zero_millimeter_if_south_of_the_south_leo_jordan_the
 NLL: 6.336
such_as_in_outcarge_of_coincination_with_mows_such_as_adler_martie_the_hilly_patt_evedhon_of_morcele_s_night_of_blood_the_tremen
t_of_eliensberg_while_an_ulav_at_esrheim_that_he_had_to_proved_left_mainied_this_label_is_in_hellenistic_separatism_the_falix_ro
 NLL: 6.805
t_orator_lemmoi_s_mother_toury_ghost_for_his_history_on_a_blaster_the_three_stallman_family_sources_including_the_film_that_a_ro
mance_nine_author_higtly_lacaded_the_second_harmour_open_source_for_which_orrie_changed_the_bluebogs_books_moy_s_athlite_s_medit
 NLL: 7.522
SFM w/o OT, ODE, NLL: 6.811, Entropy: 7.387
_became_known_as_the_shacon_valley_to_the_heaven_green_and_in_the_middle_of_the_lechneit_tracked_the_line_kej_nis_a_valley_one_p
inochules_this_was_verified_by_many_charterly_brollary_applications_including_those_which_synonymous_with_orbits_some_of_the_mas
 NLL: 6.407
cable_now_masi_had_little_to_port_from_six_eight_nine_made_hofavor_a_new_printer_of_disruption_this_platforv_would_be_faving_to_
the_current_country_but_this_need_for_saw_della_even_this_four_one_three_bit_moil_callers_did_soo_after_a_as_n_if_platform_for_t
 NLL: 6.819
nomic_ancestor_wh_meil_berg_hiarst_red_rthonstrak_utter_upon_technology_baddendin_models_on_bendrays_hypothesies_anti_aer_dynami
cs_work_have_been_intelligent_to_develop_an_european_astronomic_conifice_in_the_production_of_ten_conifices_of_develop_and_princ
 NLL: 7.479
LinearFM, ODE, NLL: 6.935, Entropy: 7.356
is_resulted_in_gawzik_college_in_the_five_season_of_feason_at_twice_the_atmosphere_is_named_after_the_list_called_him_before_inn
_s_college_at_stulpford_university_of_london_also_cambridge_the_burroughs_henrians_college_which_is_yelled_apollo_one_college_na
 NLL: 6.466
[…]
two_set_of_love_relationships_the_foundation_of_the_colfederation_homewater_to_during_the_civil_war_or_dan_brown_xian_john_zinso
 NLL: 6.935
level_mortans_already_sick_but_evade_dissolve_the_moses_of_auctional_with_deng_about_four_sekes_there_was_a_moikade_problem_to_p
eople_who_receive_signed_grief_of_culture_of_the_middle_bone_island_for_a_more_designation_of_a_kick_trade_bands_and_rangers_bom
 NLL: 7.454
MultiFlow, $T=1$, NLL: 6.728, Entropy: 7.387
er_of_the_soap_opera_by_andrew_wills_goosecat_productions_one_nine_nine_one_the_sea_monsters_of_the_late_one_nine_nine_zero_s_th
e_famous_woman_stanley_goodman_jerdre_mcnabb_out_of_zoom_movie_barry_leroy_barbara_lewis_and_brenda_punceco_aka_sney_steary_aka_
 NLL: 5.906
she_hill_obhalarnnach_eochrans_eileann_munthar_cearlomha_mhaonna__tardare_mho_mord_tore_lian_linn_mu_phaile_gael_cangallauig_lao
thuis_guilleith_leois_glion_guildh_lara_gall_innonte_tilbonne_guilecht_shuachtansert_guillaste_guatnaoic_asthache_cuichant_conai
 NLL: 6.648
lde_its_replacement_and_or_not_mist_mere_decabeod_man_and_drast_m_ek_or_ubangostrades_dialogue_or_connon_cainne_as_follows_make_
wolsey_conane_i_get_clean_to_contemplate_the_static_problem_to_reduce_it_into_perception_for_frbellist_man_jewish_views_the_othe
 NLL: 7.426
MultiFlow, $T=0.9$, NLL: 6.249, Entropy: 7.240
inism_weaver_routledge_new_york_w_g_toland_one_nine_nine_two_women_texts_gender_history_raymond_lynn_lucky_henry_pecher_one_nine
_eight_eight_the_modern_women_lyceum_new_york_harper_mead_one_nine_seven_eight_wharke_george_one_nine_nine_one_modern_history_ex
 NLL: 5.277
eight_it_deplets_the_size_of_the_starters_of_the_high_land_of_the_new_tahpu_co_bon_monte_tucals_and_quitla_land_as_elen_de_las_a
tlas_as_landous_pierce_the_torch_crack_into_steamy_places_to_eatons_hence_to_weed_out_their_blood_from_them_and_urge_them_to_hea
 NLL: 6.352
ion_of_jack_molice_now_fise_bob_springfield_slurs_boiling_funny_fruit_feed_and_chasing_rocked_duos_off_r_e_s_life_both_sides_had
_an_unwanted_and_violent_violence_and_the_rage_gained_a_recent_show_the_trial_was_the_clash_of_john_d_d_the_first_trial_at_which
 NLL: 7.074
MultiFlow, $T=0.8$, NLL: 5.552, Entropy: 6.840
ht_six_four_alexander_leroda_russian_writer_d_one_nine_one_nine_one_eight_six_seven_daniel_chase_american_author_d_one_nine_thre
e_two_one_eight_seven_eight_adolf_luroda_austrian_composer_d_one_nine_six_four_one_eight_eight_two_robert_w_keisler_american_pub
 NLL: 4.090
umadhushah_ashara_uladineshah_sahiraj_singh_one_five_seven_nine_rajnav_singh_ajmandharajaghar_singh_rohanjit_singh_one_five_eigh
t_zero_sardoyar_teachings_of_narshenhara_singh_one_five_eight_one_pranajmahr_jaharkara_jogyar_grandson_of_ujearjeer_one_five_nin
 NLL: 5.223
y_daughter_god_originates_to_come_as_a_pot_of_inspiration_all_one_of_whom_shall_witness_the_extortion_unto_you_and_will_ask_your
_neighbour_do_you_hear_and_rempskyour_lord_and_all_our_love_are_wise_tough_hope_speaked_and_thy_beautiful_names_rather_cool_when
 NLL: 6.840